The element niobium is a superconductor when cooled to temperatures below 9 K. However, the superconductivity is destroyed if the magnetic field at the surface of the metal exceeds 0.1 T. What is the maximum current that a long, straight niobium wire 3.0 mm in diameter can carry and remain a superconductor?

Please show all your work and relevant formula. Thank you.
Answer #1

Given :-

B = 0.1 T

r = d/2 = 3 mm /2 = 1.5 mm = 1.5 x 10^-3 m

uo = 4pi x 10^-7

B = uoi / 2pir

i = B*2pir / (4pi x 10^-7)

i = (0.1 x 1.5 x 10^-3) / (2 x 10^-7)

i = 750 Amp